Find all points (x, y) where the functions f(x), g(x), h(x) have the same value :
f(x) = 2^(x−5) + 3 , g(x) = 2x − 5 , h(x) = (8/x)+ 10
1
Expert's answer
2020-11-17T15:42:26-0500

First g(x) = h(x):

2x - 5 = 8/x + 10

2x - 8/x - 15 = 0

x not equal 0:

2x2 - 15x - 8 = 0

Discriminant = (-15)2 - 4 * (-8) *2 = 289 = 172

So, roots of the equation: x= (15 + 17) / (2 * 2) = 8

x2 = (15 - 17) / (2 * 2) = -0.5


Now let's analyze f(x):

1) f(8) = 2(8 - 5) + 3 = 23 + 3 = 11 equal g(8) = 2 * 8 - 5 = 11

2) f(-0.5) = 2(-0.5 - 5) + 3 = 3.022 not equal g(-0.5) = 2 * (-0.5) - 5 = -6


So the only root is x = 8.